# Exploring Robustness of LLMs to Sociodemographically-Conditioned Paraphrasing

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive performance in various NLP tasks. However, there are concerns about their reliability in different domains of linguistic variations. Many works have proposed robustness evaluation measures for local adversarial attacks, but we need globally robust models unbiased to different language styles. We take a broader approach to explore a wider range of variations across sociodemographic dimensions to perform structured reliability tests on the reasoning capacity of language models. We extend the SocialIQA dataset to create diverse paraphrased sets conditioned on sociodemographic styles. The assessment aims to provide a deeper understanding of LLMs in (a) their capability of generating demographic paraphrases with engineered prompts and (b) their reasoning capabilities in real-world, complex language scenarios. We also explore measures such as perplexity, explainability, and ATOMIC performance of paraphrases for fine-grained reliability analysis of LLMs on these sets. We find that demographic-specific paraphrasing significantly impacts the performance of language models, indicating that the subtleties of language variations remain a significant challenge. The code and dataset will be made available for reproducibility and future research [^1] https://github.com/caisa-lab.
Exploring Robustness of LLMs to
Sociodemographically-Conditioned Paraphrasing
@@ 1 Introduction
In recent years, Large Language Models (LLMs) like ChatGPT have drastically improved performance in understanding and generating human-like text, especially in tasks like question answering Tan et al. (2023). However, concerns persist regarding their robustness in understanding diverse language, particularly when presented with diverse linguistic variations inherently present in the language used by different demographic groups Marx Larre (2023); Wang et al. (2024).
@@
We introduce linguistic diversity to the LLM’s input context to address these issues and investigate their robustness to demographic paraphrases. For experimentation, we use LLAMA2-chat-13B Touvron et al. (2023), which we will refer to as LLAMA2 throughout this paper, primarily because it is open source and has shown impressive performance in NLP tasks, with comparable results, in many cases, to models like ChatGPT. For more comparison, we perform experiments with MISTRAL-instruct-7B Jiang et al. (2023), hereby referred to as MISTRAL, which is shown to outperform LLAMA2 on various NLP tasks. We make the following contributions in this paper:
  • We use LLMs for their text-generation capabilities to generate sociodemographically conditioned paraphrases. Furthermore, we identify which prompting structures are efficient based on the alignment (agreement) of the gender-oriented paraphrases with non-LLM methods.
  • We conduct robustness evalutions for LLMs in zero- and two-shot settings and contrast their performance with fine-tuned models.
  • We perform human evaluations of the paraphrases to observe correlation between QA misclassifications and paraphrase nuances.
  • We conduct fine-grained reliability tests to examine the model performance across paraphrase properties such as similarity, explainability, and atomic questions.
Our results highlight that LLMs, despite being excellent generalizers of their training corpora, struggle to reason equally efficiently with linguistic style which is more reflective of a younger demographic and which presents a more expressive language like the gender ambiguous category. Additionally, we see that in comparison to the fine-tuned models, LLMs show a similar drop in QA performance in a 0-shot setting. However, in 2-shot, we see a significant improvement compared to fine-tuned models. This is a notable finding especially because it suggests that providing LLMs with even a minimal amount of context-specific examples can dramatically improve their performance. The ability of these models to leverage just two examples to enhance their understanding and generate more accurate responses in specialized settings underlines the potential of instance-based learning.
@@ 2 Related Work
Early studies, such as those by Argamon et al. (2003), laid the foundations for linguistic attributes by demonstrating clear differences in language usage between males and females. They found that females are more likely to use pronouns, whereas males frequently use noun specifiers, highlighting an "involved" versus "informational" style in communication. Building on this, Newman et al. (2008) confirmed that not only do men and women use language differently, but these differences are consistent across different forms of communication, including social media posts and product reviews. They also suggested that these linguistic markers could predict the gender of the author with significant accuracy. Mohtasseb and Ahmed (2010) further explored this, finding that male writing style may be more consistent than female style, and that age can also influence writing style.
We explore a new approach for measuring robustness where we create demographic sets meant to adversarially test the models’ performance across different linguistic variations. We use the capabilities of the LLM itself as an agent that regenerates a given text in the requested demographic style. Here we evaluate which prompt engineering techniques are useful to obtain consistently varied paraphrases and how well the generated paraphrases align with real-world text. We also demonstrate that with minimal data augmentation, language models perform better on out-of-domain stylized text in QA tasks.
@@ 3 Methodology
Our evaluation framework consists of several components that can be seen in Figure 2. Overall, the process consists of a series of steps to analyze the robustness to paraphrased sets using the SocialIQA (SIQA) dataset. In part (A) we start by using the LLAMA2 model to create paraphrases of the SIQA dataset’s validation set, aiming to mirror different demographic linguistic styles without changing the meaning of the original context. Next, in part (B) we make sure that the paraphrases falling below a 0.8 semantic similarity threshold undergo refinement through semantic information augmented regeneration of paraphrases to ensure they preserve the original intent of the context. In (C), the alignment of these paraphrases with their intended demographic characteristics is verified using LexHub’s demographic-based lexica. Finally, we test the LLMs’ performance in question answering based on the stylized paraphrases, determining how paraphrasing affects the models’ comprehension. By comparing results with other LLMs through benchmarks like BIG-Bench Ghazal et al. (2013) (Table 5), we assess the LLMs’ capabilities against current standards, particularly in their ability to handle variations in language.
@@ 3.1 Experimental Setup
We employ LLMs, specifically LLAMA2, for paraphrasing prompts in distinct linguistic styles for the SIQA dataset as our primary corpus. Each sample from the validation set, with 1954 samples is structured in a prompt, question, and choices format, and each is independently paraphrased using the LLAMA2 model. Furthermore, for robustness evaluation on the paraphrased sets we consider two models LLAMA2 and MISTRAL. Since inference is our primary goal we use Python bindings for the lightweight llama.cpp implementation which enables LLM inference with minimal setup. For more details refer to Table 6 in the appendix.
Dataset. We construct our paraphrased QA sets using the SIQA Sap et al. (2019) dataset which comprises 38,000 multiple-choice questions focused on commonsense reasoning in social contexts. The dataset covers nine types of reasoning, such as intent, need, reaction, and effect. For our experiments, we have used the validation set with 1954 samples to get comparable benchmarks with the BIG-Bench baselines.
@@ 3.2 Paraphrase Generation
We test different prompting styles to make the language model paraphrase the given text in a particular style.
Controlled paraphrasing. The prompt consists of a basic task introduction the <<sys>>{}<</sys>> block, and a detailed instruction setup broader [INST]{}[/INST] block. We set up the LLAMA2 model for the paraphrasing task by providing the system introduction "You are an English language Expert…". Once the system is defined we provide the task information with detailed paraphrasing instructions in the INST block. Here we specify what the model precisely needs to do along with a set of rules which prevent the model from deviating from the given output format. Using this set of detailed instructions, we control the production of consistent and style-controlled paraphrases.
@@
Style-guided paraphrasing. In this method, which is an extension of the controlled paraphrasing prompt, we provide the model with detailed, explicit descriptions of the desired linguistic style traits associated with each demographic group. To guide LLAMA2 in producing paraphrases reflective of specific linguistic styles, we explore common styles associated with the given demographic group for which the paraphrasing is requested. Works such as Piersoul and Van de Velde (2023) and Preotiuc-Pietro et al. (2016) have captured distinctions in linguistic styles across different user attributes: gender, age, and occupation. Using this information, we experiment with delineating distinct linguistic attributes; for instance, for males and females [^2] Considering binary genders in this work is only one perspective and is used to highlight the differences with a sample pair. Future work can investigate how this varies for other genders., emphasizing differences in directness, emotional expression, politeness, and syntactic complexity. This approach is grounded in empirical social media text analyses and linguistic studies that outline specific stylistic tendencies Vikatos et al. (2017); Newman et al. (2008). We provide this information about the male and female style characteristics in Figure 4, with the additional instruction "Here is what you know about the linguistic style of Male and Female…".
@@
Controlling stylistic features with LexHub. LexHub is an online demographic prediction tool [^3] http://lexhub.org/wlt/lexica.html which is presented as a practical application of demographic-based lexica. By leveraging LexHub, we evaluate LLAMA2 generated paraphrases for agreement with the lexica, ensuring alignment with demographic characteristics. Sap et al. (2014) describe the creation of predictive lexica for age and gender using regression and classification models from social media data. LexHub achieved state-of-the-art accuracy in predicting age and gender over user-generated text on Facebook and Twitter.
For Age, the model directly outputs representative age for the given sample, while for Gender the score is a relative measure, where >0 implies Female and <0 implies Male. We measure the model alignment using paraphrase "agreement percentage" with the LexHub scores. First, we calculate the count of agreeing paraphrases, for Gender alignment, the following condition should hold:
$N(\texttt{lex\_score}(Par_{F})-\texttt{lex\_score}(Par_{M})>0)$
Similarly, for Age alignment, the count can be calculated by:
$N(\texttt{lex\_score}(Par_{O})-\texttt{lex\_score}(Par_{Y})>0)$
Using the above formulations, we calculate the count of paraphrase pairs where Female style paraphrase ($Par_{F}$) scores higher than the Male style paraphrase ($Par_{M}$) and Old style ($Par_{O}$) has a predicted age higher than Young style paraphrase ($Par_{Y}$). For alignment, we calculate the percentage of paraphrase pairs in Gender and Age demographic sets where the above conditions hold. It is calculated by dividing the count of agreeing pairs by the total number of samples that are paraphrased, we use 10% of randomly sampled contexts from the SIQA validation set to measure the alignment.
Using alignment as a metric to quantify the paraphrase quality, we examine which method of prompting produces better-aligned paraphrases, controlled or style-guided. Upon evaluation, it was found that with the selected hyperparameters and controlled paraphrasing, the alignment of LLAMA2 with LexHub, for Gender paraphrasing was higher with controlled paraphrasing prompt. The alignment for the controlled paraphrasing prompt was 72% while the style-guided paraphrasing presented a 64% alignment (Table 8 in the appendix). This highlights that explicitly providing the style characteristics of the groups does not necessarily help the model produce better-aligned paraphrases. Following this conclusion, we conduct our evaluation study with the paraphrased sets generated with the controlled paraphrasing prompt.
@@ 4 Robustness Evaluation
4.1 Performance on Paraphrased Sets
Figure 5: Performance with paraphrased sets contrasted with the semantic similarity of the paraphrased set with the SIQA validation set. The colored bars represent the demographic set accuracy while the gray bars show their respective cosine similarity score
The results summarized in Table 1 highlight the performance discrepancies across different demographic groups. The evaluation metrics are done on both 0-shot and 2-shot inference approaches, providing insight into the models’ ability to generalize from minimal examples. For the original test set, the LLAMA2 model demonstrates a 0-shot accuracy of 57.88%, improving to 62.42% with 2-shot inference. MISTRAL expectedly shows a superior baseline performance at 66.37% (0-shot), which improves to 71.44% (2-shot). This pattern of improvement with additional shots is consistent across demographic categories for both models, indicating that even minimal data augmentation helps in improving LLM performance.
Demographic analysis reveals nuanced performance variations. For instance, in the Male and Female categories, the models show a slight decrease in 0-shot accuracy compared to the original set, yet they improve in 2-shot, particularly notable in the Female category with LLAMA2 achieving a 63.66% (2-shot) accuracy. MISTRAL on the other hand turns out to be less robust with the female style in the 0-shot setting however shows the highest performance in the 2-shot setting with the same set. Ambiguous gender representation poses a more significant challenge, with a marked reduction in performance in both 0-shot and 2-shot inference.
Age-based disparities are evident as well. The Young demographic witnesses the most significant drop in performance for LLAMA2 in the 0-shot scenario at 54.63%, this is seen in the 2-shot setting as well, where the Young demographic shows a lower accuracy. Conversely, the Middle Age and Old groups showcase more resilient performance metrics, aligning closely with the original dataset outcomes. This resilience is particularly pronounced in the Middle Age group, where LLAMA2 achieves its highest 2-shot accuracy at 64.68%.
Performance of fine-tuned models. In this analysis, we explore the effects of demographic-specific paraphrasing on the accuracy of two fine-tuned language models, BERT-base-uncased Devlin et al. (2018) and RoBERTa-base Liu et al. (2019). Before evaluation, both models were fine-tuned from scratch on the SIQA training set for 10 epochs. Following fine-tuning, the performance of both models on various demographic paraphrases was assessed. In Table 2, we see that BERT-base-uncased started with a baseline accuracy of 56.03% on the original set. When applied to demographically paraphrased sets, there was a noticeable decline across all demographics, with the most significant drop exhibited in the Young demographic at 3.55%. The RoBERTa-base model, which began with a higher baseline accuracy of 65.76% on the original set, also showed decreased performance across the paraphrased sets, with the Ambiguous category facing the steepest decline at 4.99%.
Comparing these results to those of the foundational models, LLAMA2 and MISTRAL, a similar trend is revealed where both sets of models experience performance drops when evaluated with demographically paraphrased questions. This suggests that despite the sophistication of LLMs and better generalizing capacity, there are inherent challenges in addressing demographic-specific language that transcends the particularities of individual model architectures. Notably, the RoBERTa-base model seems to mirror the resilience of the foundational models, maintaining a relatively higher accuracy across demographic categories, much like MISTRAL.
@@ 4.1 Performance on Paraphrased Sets
Table 2: Performance of fine-tuned models (BERT and RoBERTa) on original and paraphrased SIQA sets

### v2

# Exploring Robustness of LLMs to Paraphrasing Based on Sociodemographic Factors

@@ Abstract
Despite their linguistic prowess, LLMs have been shown to be vulnerable to small input perturbations. While robustness to local adversarial changes has been studied, robustness to global modifications such as different linguistic styles remains underexplored. Therefore, we take a broader approach to explore a wider range of variations across sociodemographic dimensions. We extend the SocialIQA dataset to create diverse paraphrased sets conditioned on sociodemographic factors (age and gender). The assessment aims to provide a deeper understanding of LLMs in (a) their capability of generating demographic paraphrases with engineered prompts and (b) their capabilities in interpreting real-world, complex language scenarios. We also perform a reliability analysis of the generated paraphrases looking into linguistic diversity and perplexity as well as manual evaluation. We find that demographic-based paraphrasing significantly impacts the performance of language models, indicating that the subtleties of linguistic variation remain a significant challenge. We will make the code and dataset available for future research.
Exploring Robustness of LLMs to Paraphrasing Based on Sociodemographic Factors
@@ 1 Introduction
In recent years, Large Language Models (LLMs) like ChatGPT have drastically improved in generating human-like text, especially in tasks like question answering Tan et al. (2023). However, concerns persist regarding their robustness to diverse linguistic variations inherently present in the language used by different demographic groups Marx Larre (2023); Wang et al. (2024).
@@
We introduce linguistic diversity to the LLM’s input context to address these issues and investigate their robustness to demographic paraphrases based on age and gender. Our results highlight that LLMs, despite being impressive generalizers of their training corpora, struggle to accurately interpret the linguistic styles that are more reflective of a younger demographic and present a more expressive language like the gender-ambiguous category. Additionally, compared to the fine-tuned models, LLMs can show a similar drop in QA performance in a zero-shot setting. However, in 2-shot, we see a significant improvement compared to fine-tuned models. This is a notable finding especially because it suggests that providing LLMs with even a minimal amount of context-specific examples can dramatically improve their performance.
We make the following contributions in this paper: (1) We use LLMs for their text-generation capabilities to generate socio-demographically conditioned paraphrases, identifying which prompting structures are efficient based on the alignment (agreement) of the gender-oriented paraphrases with non-LLM methods; (2) We conduct robustness evalutions for LLMs in zero- and two-shot settings and contrast their performance with fine-tuned models; (3) We perform human evaluations of the paraphrases to investigate the correlation between QA misclassifications and paraphrase nuances; (4) We conduct fine-grained reliability tests to examine the model performance across paraphrase properties such as similarity, explainability, and atomic questions.
@@ 2 Related Work
Early studies, such as those by Argamon et al. (2003), laid the foundations for linguistic attributes by demonstrating clear differences in language usage between males and females. They found that females are more likely to use pronouns, whereas males frequently use noun specifiers, highlighting an "involved" versus "informational" style in communication. Building on this, Newman et al. (2008) confirmed that not only do men and women use language differently, but these differences are consistent across different forms of communication, including social media posts and product reviews. They also suggested that these linguistic markers could predict the gender of the author with significant accuracy. Mohtasseb and Ahmed (2010) further explored this, finding that male writing style may be more consistent than female style, and that age can also influence writing style. We explore a new approach for measuring robustness where we create demographic sets meant to adversarially test the models’ performance across different linguistic variations. We use the capabilities of the LLM itself as an agent that regenerates a given text in the requested demographic style. Here we evaluate which prompt engineering techniques are useful to obtain consistently varied paraphrases and how well the generated paraphrases align with real-world text. We also demonstrate that with minimal data augmentation, language models perform better on out-of-domain stylized text in QA tasks.
@@ 3 Methodology
Our evaluation framework consists of several components that can be seen in Figure 2. Overall, the process consists of a series of steps to analyze the robustness to paraphrased sets using the SocialIQA (SIQA) dataset. In part A, in Figure 2, we start by using the LLAMA2 model to create paraphrases of the SIQA dataset’s validation set, aiming to mirror different demographic linguistic styles without changing the meaning of the original context. Next, in part B, we make sure that the paraphrases with low semantic similarity undergo refinement through semantic information augmented regeneration of paraphrases to ensure they preserve the original intent of the context. In C, the alignment of these paraphrases with their intended demographic characteristics is verified using LexHub’s demographic-based lexica. Finally, we test the LLMs’ performance in question answering based on the stylized paraphrases, determining how paraphrasing affects the models’ comprehension.
@@ 3.1 Experimental Setup
We employ LLMs, specifically LLAMA2-chat-13B Touvron et al. (2023), which will be referred to as LLAMA2 hereafter, for paraphrasing prompts in distinct linguistic styles for the SIQA dataset as our primary corpus. Each sample from the validation set, with 1954 samples is structured in a prompt, question, and choices format, and each is independently paraphrased using the LLAMA2 model. Furthermore, for robustness evaluation on the paraphrased sets we consider two models LLAMA2 and MISTRAL-instruct-7B Jiang et al. (2023) (referred to as MISTRAL hereafter). Since inference is our primary goal we use Python bindings for the lightweight llama.cpp implementation which enables LLM inference with minimal setup. For more details refer to Table 7 in the appendix.
Dataset.
We construct our paraphrased QA sets using the SIQA Sap et al. (2019) dataset which comprises 38,000 multiple-choice questions focused on commonsense reasoning in social contexts. The dataset covers nine types of reasoning, such as intent, need, reaction, and effect. For our experiments, we have used the validation set with 1954 samples to get comparable benchmarks with the BIG-Bench Ghazal et al. (2013) baselines.
@@ 3.2 Paraphrase Generation
We test different prompting styles to make the language model paraphrase the given text in a particular style:
Controlled Paraphrasing.
The prompt consists of a basic task introduction the <<sys>>{}<</sys>> block, and a detailed instruction setup broader [INST]{}[/INST] block. We set up the LLAMA2 model for the paraphrasing task by providing the system introduction "You are an English language Expert…". Once the system is defined, we provide the task information with detailed paraphrasing instructions in the INST block. Here we specify what the model precisely needs to do along with a set of rules which prevent the model from deviating from the given output format. Using this set of detailed instructions, we control the production of consistent paraphrases.
@@
Style-guided Paraphrasing.
In this method, which is an extension of the controlled paraphrasing prompt, we provide the model with detailed, explicit descriptions of the desired linguistic style traits associated with each demographic group. To guide LLAMA2 in producing paraphrases reflective of specific linguistic styles, we explore common styles associated with the given demographic group for which the paraphrasing is requested. Works such as Piersoul and Van de Velde (2023) and Preotiuc-Pietro et al. (2016) have captured distinctions in linguistic styles across different user attributes: gender, age, and occupation. Using this information, we experiment with delineating distinct linguistic attributes; for instance, for males and females [^1] Considering binary genders in this work is only one perspective and is used to highlight the differences with a sample pair. Future work can investigate how this varies for other genders., emphasizing differences in directness, emotional expression, politeness, and syntactic complexity. This approach is grounded in empirical social media text analyses and linguistic studies that outline specific stylistic tendencies Vikatos et al. (2017); Newman et al. (2008). We provide this information about the male and female style characteristics in Figure 4, with the additional instruction "Here is what you know about the linguistic style of Male and Female…".
@@
3.3 Paraphrase Assessment
After generating the paraphrases, we assess them regarding stylistic features and correlation with perplexity:
Stylistic Features.
LexHub is an online demographic prediction tool [^2] http://lexhub.org/wlt/lexica.html which is presented as a practical application of demographic-based lexica. By leveraging LexHub, we evaluate LLAMA2 generated paraphrases for agreement with the lexica, ensuring alignment with demographic characteristics. Sap et al. (2014) describe the creation of predictive lexica for age and gender using regression and classification models from social media data. LexHub achieved state-of-the-art accuracy in predicting age and gender over user-generated text on Facebook and Twitter.
For Age, the model directly outputs representative age for the given sample, while for Gender the score is a relative measure, where >0 implies Female and <0 implies Male. We measure the model alignment using paraphrase agreement percentage with the LexHub scores. First, we calculate the count of agreeing paraphrases; for Gender alignment, the following condition should hold:
$N(\texttt{lex\_score}(Par_{F})-\texttt{lex\_score}(Par_{M})>0)$
Similarly, for Age alignment, the count can be calculated by:
$N(\texttt{lex\_score}(Par_{O})-\texttt{lex\_score}(Par_{Y})>0)$
Using the above formulations, we calculate the count of paraphrase pairs where Female style paraphrase ($Par_{F}$) scores higher than the Male style paraphrase ($Par_{M}$) and Old style ($Par_{O}$) has a predicted age higher than Young style paraphrase ($Par_{Y}$). For alignment, we calculate the percentage of paraphrase pairs in Gender and Age demographic sets where the above conditions hold. It is calculated by dividing the count of agreeing pairs by the total number of samples that are paraphrased. We use 10% of randomly sampled contexts from the SIQA validation set to measure the alignment.
Prompt style LexHub Alignment Cumulative scores
Ambig Female Male
Controlled 72% 1.33 235.90 -248.36
Style-guided 64% 12.49 112.67 -86.97
Table 1: Paraphrase alignment with LexHub for controlled vs. style-guided paraphrasing with LLAMA2. The cumulative scores show the raw aggregated scores for the paraphrased samples. Ambig is short for ambiguous.
Using alignment as a metric to quantify the paraphrase quality, we examine which method of prompting produces better-aligned paraphrases, controlled or style-guided. Upon evaluation, we found that the alignment of LLAMA2 with LexHub using controlled paraphrasing for Gender paraphrasing was higher (72%) than the style-guided paraphrasing (64%) (Table 1). This highlights that explicitly providing the style characteristics of the groups does not necessarily help the model produce better-aligned paraphrases. Following this conclusion, we conduct our evaluation study with the paraphrased sets generated with the controlled paraphrasing prompt.
Correlation with Perplexity.
Perplexity is a measure of a model’s uncertainty in predicting the next word in a sequence. In the context of linguistic variation, given that the content of the sentence is the same, perplexity can also be interpreted as how comfortable the model is with predicting that style. The perplexity scores reported in Table 2 for LLAMA2 and GPT-2 across different demographic categories reveal significant insights into model behavior. Notably, we see an inverse relationship between perplexity and model performance; as perplexity increases, we observe a corresponding decrease in model accuracy. Figure 5 shows this correlation. We consider GPT-2 as an unbiased third model to calculate the set perplexity because the language is generated by the LLM itself. For GPT-2, results with the 0-shot setup demonstrate an inverse correlation between perplexity and model accuracy and higher perplexity scores directly translate to lower accuracy. This is also evident, albeit not so strongly, in the 2-shot setting. This possibly indicates that when the model is allowed to learn the language nuances in a low data setup, language complexity (perplexity) plays a smaller role.
@@ 4 Robustness Evaluation
4.1 Decoder Models
@@
Table 3: Performance of encoder models (BERT and RoBERTa) on original and paraphrased SIQA sets
The results summarized in Table 2 highlight the performance discrepancies across different demographic groups. The evaluation metrics are done on both 0-shot and 2-shot inference approaches, providing insight into the models’ ability to generalize from minimal examples. For the original test set, the LLAMA2 model demonstrates a 0-shot accuracy of 57.88%, improving to 62.42% with 2-shot inference. MISTRAL expectedly shows a superior baseline performance at 66.37% (0-shot), which improves to 71.44% (2-shot). This pattern of improvement with additional shots is consistent across demographic categories for both models, indicating that even minimal set of examples helps in improving LLM performance.
Demographic analysis reveals nuanced performance variations. For instance, in the Male and Female categories, the models show a slight decrease in 0-shot accuracy compared to the original set, yet they improve in 2-shot, particularly notable in the Female category with LLAMA2 achieving a 63.66% (2-shot) accuracy. MISTRAL, on the other hand, turns out to be less robust with the female style in the 0-shot setting while showing the highest performance in the 2-shot setting with the same set. Ambiguous gender representation poses a more significant challenge, with a marked reduction in performance in both 0-shot and 2-shot inference.
Age-based disparities are evident as well. The Young demographic witnesses the most significant drop in performance for LLAMA2 in the 0-shot scenario at 54.63%. This is seen in the 2-shot setting as well, where the Young demographic shows a lower accuracy. Conversely, the Middle Age and Old groups showcase more resilient performance metrics, aligning closely with the original dataset outcomes. This resilience is particularly pronounced in the Middle Age group, where LLAMA2 achieves its highest 2-shot accuracy at 64.68%.
4.2 Encoder Models
In this analysis, we explore the effects of demographic-specific paraphrasing on the accuracy of two encoder models, BERT-base-uncased Devlin et al. (2018) and RoBERTa-base Liu et al. (2019). Before evaluation, both models were fine-tuned on the SIQA training set for 10 epochs. Following fine-tuning, the performance of both models on various demographic paraphrases was assessed. In Table 3, we see that BERT-base-uncased started with a baseline accuracy of 56.03% on the original set. When applied to demographically paraphrased sets, there was a noticeable decline across all demographics, with the most significant drop exhibited in the Young demographic at 3.55%. The RoBERTa-base model, which began with a higher baseline accuracy of 65.76% on the original set, also showed decreased performance across the paraphrased sets, with the Ambiguous category facing the steepest decline at 4.99%.
Comparing these results to those of the decoder models, LLAMA2 and MISTRAL, a similar trend is revealed where both sets of models experience performance drops when evaluated with demographically paraphrased questions. This suggests that despite the sophistication of LLMs and better generalizing capacity, there are inherent challenges in addressing demographic-specific language that transcends the particularities of individual model architectures.